I need help please.
A foot tall tree casts a shadow that is one foot shorter than the distance from the tip of the shadow to the top of the tree. Find the length of shadow.
(a) Let x represent the length of the shadow. Write an expression for the diagonal distance from the tip of the shadow to the top of the tree.
(B) What is the length of the shadow? (Hint use the pythagorean formula)

Answer by rothauserc(4718)

a) diagonal distance is x+1
b) 1 + x^2 = (x+1)^2
1 + x^2 = x^2 +2x +1
2x = 0
x = 0 implies the length of the shadow is zero